A series RLC circuit consists of a 20 Ω resistor, a 51 μF capacitor, and a 25 mH inductor. If the source frequency is 50 Hz, and the circuit current is 350 mA, what is the applied voltage?

1
Expert's answer
2022-01-25T02:36:02-0500

Vr= 20*0.35=7 volt

XL=2*3.14*f*L= 2*3.14**50*25* 10-3=7.85 ohm

Xc= 1/2*3.14*f*C= 1/2*3.14*50*51*10-6=62.44 ohm

Xc-XL=54.6 ohm

Vx=I*(Xc-XL)=19.1 Volt

Vs= (Vx2+Vr2)1/2= 20.35 Volt